SOLVE NUMERICALLY USING EULERS METHOD

The motion of a damped spring-mass system (Fig. P25.16)  is described by the following ordinary differential equation: m d2x dt2 1 c dx dt 1 kx 5 0 where x 5 displacement from equilibrium position (m), t 5 time  (s), m 5 20-kg mass, and c 5 the damping coeffi cient (N ? s/m).  The damping coeffi cient c takes on three values of 5 (underdamped), 40 (critically damped), and 200 (overdamped). The  spring constant k 5 20 N/m. The initial velocity is zero, and the  initial displacement x 5 1 m. Solve this equation using a numerical  method over the time period 0 # t # 15 s. Plot the displacement  versus time for each of the three values of the damping coeffi cient  on the same curve.


SOLVE NUMERICALLY USING EULERS METHOD
Answer #1

Solution:

It is required solve the equation of motion of spring-mass system using numerical method for the three damping conditions andTo solve the required problem, write the following code in the editor window of the Matlab to implement RK4 method and save i16 - 17 18 - 19 20 21 22 if nargin<4, error(at least 4 input arguments required), end $checking number of input arguments ielse 32 33 34 35 - 36 37 - 38 39 - I 40 41 42 t = tspan; end tt = ti; y(1,:) = y0; np = 1; tp (np) = tt; yp (np, :) = y(1,:);51 - 52 53 - 54 55 - 56 - k4 = dydt (tt+hh, vend,varargin{:}); phi = (kl+2* (k2+k3) +k4)/6; y(i+1, :) y(i,:) + phi*hh; tt =1 2 - 3- The codes are explained in the codes itself. The sentences starting with % is the explanation of the codes written jIt is to be noted that the above function calls the function dydt, which is equation of motion for the spring mass system. WrNow, call the function prob22150) by typing the following command in the command window of the Matlab: fx >> prob2215 The outThe plot generated as an output is shown below: Displacements for mass-spring system 0.8 00 0.6 0.4 0.2 x (m) 0 -0.2 -0.4 -0.


answered by: ANURANJAN SARSAM